What are the critical values of for area in two tails when the sample size is 29 with confidence level of 98%?




1
Expert's answer
2021-06-07T19:31:34-0400

The test statistic follows the t-distribution with "n-1=29-1=28" 28 degrees of freedom.

"\alpha=0.02" for 98% level

Critical value: ±2.4671